OpenStudy (anonymous):

3/4% x75 explain/check answer. 3/4 *100 x 75

OpenStudy (anonymous):

A. .255 B .5625 C. 2.55 D 5.625

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

I think you've got it :) (3/4) x (1/100) x 75 = (3/400) x 75 = Answer is D.

OpenStudy (anonymous):

i cant seem to get any of the choices

OpenStudy (danjs):

three fourths percent of 75?

OpenStudy (danjs):

\[\frac{ 0.75 }{ 100 } *75\]

OpenStudy (anonymous):

keep in mind that's 3/4% so it's 3/4 out of a 100

OpenStudy (anonymous):

what about 5/10 x 55? 5/100 x 55= 2.75 right???

OpenStudy (anonymous):

did you mean (5/10)% or just (5/10)?

OpenStudy (anonymous):

oh percent

OpenStudy (anonymous):

ok. then \[\frac{ \frac{ 5 }{ 10 } }{ 100 } \times 55\] = \[\frac{ 5 }{ 10 } \times \frac{ 1 }{ 100 } \times 55\]

OpenStudy (anonymous):

your answer should have been (5/1000) x 55 = 0.275 but you were close

OpenStudy (anonymous):

oh i was lol